OpenStudy (anonymous):

What is the length of a pendulum that swings back and forth with a frequency of 2.5 Hz?

mathslover (mathslover):

Use the formula : \(\large{\boxed{\sf{T = 2\pi \sqrt{\cfrac{L}{g}}}}}\) T is the time period of the Pendulum L is the Length of a pendulum g is the acceleration due to gravity which can be taken as \(\sf{9.8 ms^{-2}}\) \(\pi\) is constant which is equal to 3.14 *Time period is the reciprocal of Frequency.* \(T = \cfrac{1}{\nu}\) You are given with Frequency (\(\nu\)) = 2.5 Hz Find T and then put the values in the first formula.

mathslover (mathslover):

Then you can find the Length of a pendulum !